OpenStudy (anonymous):

Simplify (a^2b^-3)^4

OpenStudy (anonymous):

Do you mean b^-3?

OpenStudy (campbell_st):

the power of r opertaes on both terms so it can be written as \[(a^2)^4 \times (b^{-3})^4\] and now you need the power of a power law \[(x^a)^b = x^{a \times b}\] which requires you to multiply the powers. hope it helps

OpenStudy (triciaal):

using the information from @campbell_st you should have this